# A HHO formulation for variable density incompressible flows where the density is purely advected

### 04 Mar 2026 v3

@@ 3.7 Stability
where, in the first step, we introduced the definition of the gradient reconstruction operator, and in the second step, we used the fact that the element velocity $\bm{u}_{T}$ is divergence-free and H-div conforming, see Sec. 3.5, together with the single-valuedness of $p_{\partial T}$ on internal faces. The non-dissipation property follows, since, in case of homogeneous Dirichlet BCs, $\bm{u}_{T}\cdot\bm{n}_{TF}=\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF})=0$. Accordingly, we remark that all the terms on the right hand side of (19)-(20) but $\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}$ are non-negative. As demonstrated in [13, Ch. 5.1.6], see also [11, Remark 2.2], the HHO discretization of the Laplace operator introduced in (5) is coercive i.e., for all $T\in\mathcal{T}_{h}$ and all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, it holds
@@ 4 Numerical results
We remark that Newton’s method is solved up to machine precision in order to guarantee exact conservation of volume at the discrete level, accordingly the linear solver stopping criterion does not influence the numerical results. In order to further improve the performance of the iterative linear solver required for the solution of each Newton step we rely on the $p$-multilevel preconditioner strategy proposed in [7]. In particular, we apply one sweep of $p$-multigrid V-cycle as a preconditioner for the FGMRES (flexible GMRES) iteration. As a smoothing strategy within the multigrid V-cycle preconditioner we rely on a few iterations of GMRES. All applications of prolongation and restriction operators involved in the V-cycle iteration are performed matrix-free, that is, without assembling the global sparse matrices associated to the operators, see [7] for details.
@@ 5 Conclusions
Further research efforts are required to target physically relevant multi-phase flows, in particular, future works will consider the extension of the HHO-ESDIRK formulation to the Cahn-Hilliard-Navier-Stokes (CHNS) system.

### 02 May 2026 v4

@@ 3.7 Stability
where, in the first step, we introduced the definition of the gradient reconstruction operator, and in the second step, we used the fact that the element velocity $\bm{u}_{T}$ is divergence-free and H-div conforming, see Sec. 3.5, together with the single-valuedness of $p_{\partial T}$ on internal faces. The non-dissipation property follows, since, in case of homogeneous Dirichlet BCs, $\bm{u}_{T}\cdot\bm{n}_{TF}=\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF})=0$. Accordingly, we remark that all the terms on the right hand side of (27)-(28) but $\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}$ are non-negative. As demonstrated in [17, Ch. 5.1.6], see also [15, Remark 2.2], the HHO discretization of the Laplace operator introduced in (5) is coercive i.e., for all $T\in\mathcal{T}_{h}$ and all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, it holds
@@ 4 Numerical results
We remark that Newton’s method is solved up to machine precision in order to guarantee exact conservation of volume at the discrete level, accordingly the linear solver stopping criterion does not influence the numerical results. In order to further improve the performance of the iterative linear solver required for the solution of each Newton step we rely on the $p$-multilevel preconditioner strategy proposed in [8]. In particular, we apply one sweep of $p$-multigrid V-cycle as a preconditioner for the FGMRES (flexible GMRES) iteration. As a smoothing strategy within the multigrid V-cycle preconditioner we rely on a few iterations of GMRES. All applications of prolongation and restriction operators involved in the V-cycle iteration are performed matrix-free, that is, without assembling the global sparse matrices associated to the operators, see [8] for details. Static condensation is performed cell-by-cell at matrix assembly, meaning that only statically condensed global matrices are stored in memory. It is worth to mention that the code can be executed in parallel based on a distributed memory paradigm and relies on the iterative solvers and preconditioners provided by the PETSc library [1]. From the computational cost viewpoint, as is often the case when considering fully implicit time-marching strategies for incompressible fluid flow computations, the computational expense associated with linear systems solutions dominates over matrix assembly CPU-times. In practice, the ESDIRK-HHO formulation coupled with the $p$-multigrid solution strategy is best efficient when considering higher-order polynomials and coarse meshes.
@@ 5 Conclusions
Further research efforts are needed to address physically relevant multi-phase flows, in particular, future works will consider extending the present HHO-ESDIRK formulation to the Cahn-Hilliard-Navier-Stokes (CHNS) system. HHO formulations that include heterogenous viscosity coefficients and strain-rate related viscous stresses are already available in literature, see, e.g.,  [9, 12], alongside, HHO methods have already been succesfully applied to the Cahn-Hilliard problem, see e.g.,  [14]. We plan to investigate up to which extent the intrinsic interface thickness control provided by the Cahn-Hilliard model will help in mitigating the onset of spurious oscillations when higher-order ($k>0$) formulations are considered.